### v2

@@ Abstract
Transforming CO2 into methanol represents a crucial step towards closing the carbon cycle, with thermoreduction technology nearing industrial application. However, obtaining high methanol yields and ensuring the stability of heterocatalysts remain significant challenges. Herein, we present a sophisticated computational framework to accelerate the discovery of novel thermal heterogeneous catalysts, using machine-learned force fields. We propose a new catalytic descriptor, termed adsorption energy distribution, that aggregates the binding energies for different catalyst facets, binding sites, and adsorbates. The descriptor is versatile and can easily be adjusted to a specific reaction through careful choice of the key-step reactants and reaction intermediates. By applying unsupervised machine learning and statistical analysis to a dataset comprising nearly 160 metallic alloys, we offer a powerful tool for catalyst discovery. Finally, we propose new promising candidate materials such as ZnRh and ZnPt3, which to our knowledge, have not yet been tested, and discuss their possible advantage in terms of stability.
@@ 2 Results and Discussion
To identify the most crucial adsorbates for AEDs calculations, we perused the existing literature. An experimental investigation by Amman et al. [30] highlighted the presence of surface-bound radicals such as *H (hydrogen atom), *OH (hydroxy group), *OCHO (formate), and *OCH3 (methoxy) as essential reaction intermediates in the thermocatalytic reduction of CO2 to methanol. Based on these findings, we selected them for our AEDs calculations. Please note that the notation for formate can vary across the literature, e.g., *HCOO [30, 39] and HCOO* [40]. With the help of fairchem repository tools by OCP [41], we created surfaces with their Miller index $\in\{-2,-1,...,2\}$ and calculated their total energy using OCP MLFF. If we encountered multiple cuts for the same facet, we selected the one with lowest energy for further calculations. Then we engineered surface-adsorbate configurations for the most stable surface terminations across all facets within our defined Miller index range for the materials, as described in section 3.2, and optimized these configurations using the OCP MLFF. During this process, we discovered that six materials exhibited so large surface-adsorbate supercells, that their calculations were infeasible on available GPU resources, even with the effective OCP MLFF. Consequently, they were excluded from our study.
@@
Figure 4: Hierarchical clustering of the materials, based on the Wasserstein Distances: The graph show that all 159 materials were assembled into 19 clusters, based on the similarity between of the AEDs, with the exception of potassium, which is dissimilar to nearby materials and forms a single material, non-numbered cluster. This can be seen in detail in the insets, where also materials in clusters 8, 9 and 10 are can be seen. The cluster 10 contains several alloys, which are part of known high-yield catalysts as well as new potentially active materials.
For a clustering threshold corresponding to a Wasserstein distance of 0.025, we arrive at a total of 19 distinct clusters, with potassium (K) forming its own, isolated, unnumbered cluster. The separation between clusters 11 to 19 and clusters 1 to 10 is considerable. The distinguishing feature is the broadness of the AEDs. The distributions in clusters 11 to 19 are noticeably broader than in clusters 1 to 10. Representative examples are depicted in Fig. 3(a). The four materials at the bottom of the figure (Y3Zn11, V6Ga5, Y3In5, V4Zn5) pertain to cluster 18, whereas the rest belong to clusters 8 to 10. Further details are available in Fig. S1 of the Supplementary Information, presenting the clustering of all considered materials. AEDs exhibit variability across distinct clusters (1 – 10) but show remarkable similarity within each individual cluster. For example, the AEDs for Ni, CuZn, NiZn, and ZnRh illustrated in Fig. 3(d-g) belong to the same cluster.
@@
Upon closer inspection, ZnRh and ZnPt3 stand out as new candidates. They are part of cluster 10, which also includes Ga2Cu, NiZn, InPt3, Ni and mainly CuZn, but have not been tested for CO2 to methanol conversion. While the exact material composition of the Cu and ZnO-based catalysts during the proceeding reaction is still debated [46, 47, 48], some studies suggest that the formation of a Cu-Zn alloy enhances the activity. This Cu-Zn alloy, which is part of cluster 10, is believed to contribute to increased activity [47, 49, 30]. Similarly, NiZn has also been identified as an effective CO2 catalyst [45]. Catalysts such as Cu/Ga [50] and Pt/In2O3 [51], known for their high methanol yield, may include Ga2Cu and InPt3 alloys, respectively. Finally, Ni is often part of catalysts for CO2 transformation to methane [52, 53]. The strong catalytic activity of Ga2Cu, NiZn, InPt3, Ni and mainly CuZn in this cluster suggest that also ZnRh and ZnPt3 should have a high activity.
@@
AEDs could serve as a descriptor of activity, however, the vast number of parameters (at least 388 bins in the distribution) makes it challenging to analyze them manually. To further our insight into the generated data, we conducted a statistical analysis of our AEDs (SAAEDs) that facilitates comparison with previous adsorption energy-based studies. An example can be seen in Table 2, where we present the minimum adsorption energies for a subset of materials featured in Fig 3(a).
@@
Both catalyst descriptors are tailored to perform extensive searches for catalytically active materials candidates. They do, however, not include effects of the support, additives, preparation procedures and operando states that could change the morphology of the catalyst (e.g., nanoparticle sizes or areas of different facets). Moreover, our AED descriptor does not take the facet area into account, and is therefore insensitive to morphology changes of the catalysts under reaction conditions. In principle, a Wulff construction could introduce better facet information. However, it also cannot account for support effects, additives or preparation conditions. Incorporating general descriptors (e.g., Magpie) for support materials could provide additional information to decide which active material-support combination should be investigated further in experimental testing
@@ 3.5 Unsupervised Learning
Acknowledgements
The authors would like to thank Annukka Santasalo-Aarnio and Arpad Toldy for fruitful discussions. O.K. and P.P. express their gratitude to Kirby Broderick, Adeesh Koluru, Brook Wander, John Kitchin and other Kitchin research group members at Carnegie Mellon University and Zachary Ulissi at Meta, for their help with the OCP models. This project received funding from the European Union – NextGenerationEU instrument and the Research Council of Finland’s AICon project (grant number no. 348179). The authors gratefully acknowledge CSC – IT Center for Science, Finland, and the Aalto Science-IT project for the generous computational resources.
Declarations
The authors declare no conflicts of interest. The data published in this study can be found in [43]. P.P. and O.K. created the workflow and proceeded the calculation. P.R. supervised the work. All authors contributed to the manuscript.

### v3

@@ Abstract
Transforming CO2 into methanol represents a crucial step towards closing the carbon cycle, with thermoreduction technology nearing industrial application. However, obtaining high methanol yields and ensuring the stability of heterocatalysts remain significant challenges. Herein, we present a sophisticated computational framework to accelerate the discovery of thermal heterogeneous catalysts, using machine-learned force fields. We propose a new catalytic descriptor, termed adsorption energy distribution, that aggregates the binding energies for different catalyst facets, binding sites, and adsorbates. The descriptor is versatile and can be adjusted to a specific reaction through careful choice of the key-step reactants and reaction intermediates. By applying unsupervised machine learning and statistical analysis to a dataset comprising nearly 160 metallic alloys, we offer a powerful tool for catalyst discovery. We propose new promising candidates such as ZnRh and ZnPt3, which to our knowledge, have not yet been tested, and discuss their possible advantage in terms of stability.
@@ 2 Results and Discussion
To identify the most crucial adsorbates for AEDs calculations, we perused the existing literature. An experimental investigation by Amman et al. [30] highlighted the presence of surface-bound radicals such as *H (hydrogen atom), *OH (hydroxy group), *OCHO (formate), and *OCH3 (methoxy) as essential reaction intermediates in the thermocatalytic reduction of CO2 to methanol. Based on these findings, we selected these adsorbates for our AEDs calculations. Please note that the notation for formate can vary across the literature, e.g., *HCOO [30, 39] and HCOO* [40]. With the help of fairchem repository tools by OCP [41], we created surfaces with their Miller index $\in\{-2,-1,...,2\}$ and calculated their total energy using OCP MLFF. If we encountered multiple cuts for the same facet, we selected the one with lowest energy for further calculations. Then we engineered surface-adsorbate configurations for the most stable surface terminations across all facets within our defined Miller index range for the materials, as described in section 3.2, and optimized these configurations using the OCP MLFF. During this process, we discovered that six materials exhibited so large surface-adsorbate supercells, that their calculations were infeasible on available GPU resources, even with the effective OCP MLFF. Consequently, they were excluded from our study.
@@
For a clustering threshold corresponding to a Wasserstein distance of $2.5\times 10^{-3}$, we arrive at a total of 19 distinct clusters, with potassium (K) forming its own, isolated, unnumbered cluster. The separation between clusters 11 to 19 and clusters 1 to 10 is considerable. The distinguishing feature is the broadness of the AEDs. The distributions in clusters 11 to 19 are noticeably broader than in clusters 1 to 10. Representative examples are depicted in Fig. 3(a). The four materials at the bottom of the figure (Y3Zn11, V6Ga5, Y3In5, V4Zn5) pertain to cluster 18, whereas the rest belong to clusters 8 to 10. Further details are available in Fig. S1 of the Supplementary Information, presenting the clustering of all considered materials. AEDs exhibit variability across distinct clusters (1 – 10) but show remarkable similarity within each individual cluster. For example, the AEDs for Ni, CuZn, NiZn, and ZnRh illustrated in Fig. 3(d-g) belong to the same cluster.
Figure 4: Hierarchical clustering of the materials, based on the Wasserstein distances: The graph shows that all 159 materials were assembled into 19 clusters, based on the similarity between of the AEDs, with the exception of potassium, which is dissimilar to nearby materials and forms a single, non-numbered cluster. This can be seen in detail in the insets, where materials in clusters 8, 9 and 10 are shown. The cluster 10 contains several alloys, which are part of known high-yield catalysts as well as new potentially active materials.
@@
Upon closer inspection, ZnRh and ZnPt3 stand out as new candidates. They are part of cluster 10, which also includes Ga2Cu, NiZn, InPt3, Ni and mainly CuZn, but have not been tested for CO2 to methanol conversion. While the exact material composition of the Cu and ZnO-based catalysts during the proceeding reaction is still debated [46, 47, 48], some studies suggest that the formation of a Cu-Zn alloy enhances the activity. This Cu-Zn alloy, which is part of cluster 10, is believed to contribute to increased activity [47, 49, 30]. Similarly, NiZn has also been identified as an effective CO2 catalyst [45]. Catalysts such as Cu/Ga/ZnO [50], Cu/Ga/SiO2 [51] and Pt/In2O3 [52], known for their high methanol yield, may include Ga2Cu and InPt3 alloys, respectively. Finally, Ni is often part of catalysts for CO2 transformation to methane [53, 54]. The strong catalytic activity of Ga2Cu, NiZn, InPt3, Ni and mainly CuZn in this cluster suggest that also ZnRh and ZnPt3 should have a high activity.
@@
AEDs could serve as a descriptor of activity, however, the vast number of parameters (at least 388 bins in the distribution) makes it challenging to analyze them manually. To further our insight into the generated data, we conducted a statistical analysis of AEDs (SAAEDs) that facilitates comparison with previous adsorption energy-based studies. An example can be seen in Table 2, where we present the minimum adsorption energies for a subset of materials featured in Fig 3(a).
@@
Both catalyst descriptors are tailored to perform extensive searches for catalytically active candidates. They do not, however, include effects of the support, additives, preparation procedures and operando states that could change the morphology of the catalyst (e.g., nanoparticle sizes or areas of different facets). Our proposed AED descriptor does not take the facet area into account, and is therefore insensitive to morphology changes of the catalysts under reaction conditions. In principle, a Wulff construction could introduce better facet information. However, it also cannot account for support effects, additives or preparation conditions.
Our methodology facilitates high-throughput screening of metallic catalyst candidates. At present, effects of co-operating oxides such as ZnO, In2O3, and ZrO2 that have been observed experimentally, are not considered. Such oxides affect the electronic structure and adsorption energy landscape at the metal-oxide interface [56, 57, 55, 58, 46] and should be included in future versions of our descriptor. For instance, incorporating general descriptors (e.g., Magpie) for co-catalysts and support materials could provide additional information to decide which active material-support combination should be investigated further in experimental testing.
Additionally, the choice of adsorbates can also influence the effectiveness of the AED descriptor. Our study focuses on four most relevant intermediates, observed on Cu(211) surface [30]. Studies on different materials, such as Ni-ZrO2 [58, 59], suggest that other intermediates or by-products like CO could play an important role in the hydrogenation mechanism. Thus, further investigations could extend the set of adsorbates to better capture various reaction paths and therefore material-specific activity.
@@ 3.5 Unsupervised Learning
Data Availability: The data published in this study can be found in [43].
[…]
Acknowledgements The authors would like to thank Annukka Santasalo-Aarnio and Arpad Toldy for fruitful discussions. O.K. and P.P. express their gratitude to Kirby Broderick, Adeesh Koluru, Brook Wander, John Kitchin and other Kitchin research group members at Carnegie Mellon University and Zachary Ulissi at Meta, for their help with the OCP models. This project received funding from the European Union – NextGenerationEU instrument and the Research Council of Finland’s AICon project (grant number no. 348179). The authors gratefully acknowledge CSC – IT Center for Science, Finland, and the Aalto Science-IT project for the generous computational resources.
[…]
Author contributions P.P. and O.K. created the workflow and proceeded the calculations. P.R. supervised the work. All authors contributed to the manuscript.
[…]
Competing Interests The authors declare no conflicts of interest.